OpenStudy (anonymous):

(INSTANT MEDAL & FAN) How many roots does y = x^2 - 3x + 4 have? um...is it 3?

jimthompson5910 (jim_thompson5910):

The degree is _____

jimthompson5910 (jim_thompson5910):

If you're not sure what the degree is, it's the largest exponent

OpenStudy (anonymous):

2

OpenStudy (anonymous):

the degree is, your cute, lol

OpenStudy (anonymous):

thanks taylor :)

OpenStudy (anonymous):

@kevi23 thanks :)

OpenStudy (anonymous):

ur, welcome @ChinaGurl

OpenStudy (anonymous):

Set equal to 0: x^2 - 3x + 4 = 0 Now find a, b such that a + b = -3 and ab = 4 Let a = -4, b = 1: (x - 4)(x + 1) = 0 So either (x-4) = 0 or (x+1) = 0... what situations can this happen in? Either x = 4 or x = -1. Nothing else will cause one of those terms to equal zero. That's why it has 2 roots.